1. Sam invested some money at 7% and then invested $4,000 more than twice this amount at 9%. His total annual income from the two investments was $2,860. How much was invested at 9%?


2. How many liters of a 10% alcohol soultion must be mixed with 30 liters of a 40% solution to get a 20% solution?

Answer by stanbon(75887)

Sam invested some money at 7% and then invested $4,000 more than twice this amount at 9%. His total annual income from the two investments was $2,860. How much was invested at 9%?
---
0.07x + 0.09(2x+4000) = 2850
Multiply thru by 100 to get:
7x + 18x + 36000 = 285000
25x = 249000
x = $9960 (amt invested at 7%)
2x+4000 = 2*9960+4000 = $23,920 (amt. invested at 9%
======================================================

2. How many liters of a 10% alcohol solution must be mixed with 30 liters of a 40% solution to get a 20% solution?
---
alcohol + alcohol = alcohol
0.10x + 0.40*30 = 0.20(x+30)
Multiply thru by 100 to get:
10x + 40*30 = 20x + 20*30
10x = 20*30
x = 60 liters (amt. of 10% solution needed for the mixture)
===========================
Cheers,
Stan H.
===========================